# CAN QUANTUM THEORY BE UNDERPINNED BY A NON-LOCAL HIDDEN VARIABLE THEORY ?

### v1

@@ 5.3 Quantum Theory Test
and using the QTHY result for $P(\alpha_{k},\beta_{l}|\Omega_{Aa},\Omega_{Bb},\rho)_{Q}$ in the case of the maximally entangled state we obtain
$S$ $=$ $\frac{1}{2d^{3}}\;\mathop{\displaystyle\sum}\limits_{jl}\frac{1}{\sin^{2}(\frac{\pi}{d}(j-l+\theta_{1}+\phi_{1}))}\;\delta_{j,l+\Delta_{11}}+\frac{1}{2d^{3}}\;\mathop{\displaystyle\sum}\limits_{kl}\frac{1}{\sin^{2}(\frac{\pi}{d}(k-l+\theta_{2}+\phi_{1}))}\;\delta_{l,k+\Delta_{12}}$ (42)
$+\frac{1}{2d^{3}}\;\mathop{\displaystyle\sum}\limits_{km}\frac{1}{\sin^{2}(\frac{\pi}{d}(k-m+\theta_{2}+\phi_{2}))}\;\delta_{k,m+\Delta_{22}}+\frac{1}{2d^{3}}\;\mathop{\displaystyle\sum}\limits_{jm}\frac{1}{\sin^{2}(\frac{\pi}{d}(j-m+\theta_{1}+\phi_{2}))}\;\delta_{m,j+\Delta_{21}}$
@@
Note that adding an integer multiple of $d$ to the $\Delta_{ij}$ does not alter the result.
We could choose $\Delta_{ij}$ in the form $\Delta_{ij}=d\times\Delta{}_{ij}^{\ast}$ where the $\Delta_{ij}^{\ast}$ are all integers. In this case, using $\sin^{2}(\pi\Delta_{ij}^{\ast}+\eta)=(\sin(\pi\Delta_{ij}^{\ast})\cos(\eta)+\cos(\pi\Delta_{ij}^{\ast})\sin(\eta))^{2}=\sin^{2}(\eta)$ we finally obtain
@@
This last result shows that in quantum theory the value of $S$ is independent of the $\Delta_{ij}^{\ast}$. It only depends on $d$. The NLHVT Bell inequalities for $S$ do of course depend on the $\Delta_{ij}^{\ast}$.
For the choice $d=10$ we find that $S\approx 2.52$, so this exceeds $S\leq 2$ and $S\leq 1$ though not $S\leq 4$ and $S\leq 3$ that could occur in NLHVT depending on the requirements for these cases being met. For the choice $(\Delta_{11}^{\ast},\Delta_{12}^{\ast},\Delta_{22}^{\ast},\Delta_{21}^{\ast})=(1,-2,1,-2)$ it is easily confirmed that the requirements for $S\leq 4$ are not met. However the requirements for $S\leq 3$ are met, as are the reqirements for $S\leq 1$, though not the requirements for $S\leq 2$. Other choices are also be made. For example with $(\Delta_{11}^{\ast},\Delta_{12}^{\ast},\Delta_{22}^{\ast},\Delta_{21}^{\ast})=(1,-3,1,-3)$ it is easily confirmed that the requirements for $S\leq 4$ are not met. However the requirements for $S\leq 3,S\leq 2$ and $S\leq 1$ are all met. Thus we have found cases for $d=10$ where the quantum theory prediction that $S=2.52$ is inconsistent with the NLHVT Bell inequality prediction of $S\leq 2.$
We could also choose $(\Delta_{11},\Delta_{12},\Delta_{22},\Delta_{21})=(0,1,0,0)$ as in the original CGLMP Bell inequality. In this case the Bell inequality predicts $S\leq 3$. For this choice the expression for $S$ is given by
@@
For the case of $d=2$ this gives $S=3.41$. For the choice $(\Delta_{11},\Delta_{12},\Delta_{22},\Delta_{21})=(0,1,0,0)$ the requirements for $S\leq 4$ are not met, but those for $S\leq 3$ are met,.The requirements for $S\leq 4$, $S\leq 2$ and $S\leq 1$ are not met. However NLHVT is inconsistent with QTHY in this choice of $(\Delta_{11},\Delta_{12},\Delta_{22},\Delta_{21})$ as it was for LHVT where in both  types of HVT the inequality $S\leq 3$ is violated in Q THY.
@@ 6 SUMMARY AND CONCLUSION
A brief review of the origins of local hidden variable theory, Bell local states and Bell inequalities has been presented. The classification of Bell local states into three categorise based on the local hidden quantum state approach of Wiseman has been explained. We outline how the Bell inequalities provide tests for whether local hidden variable theory is in conflict with quantum theory and also which categories of Bell local states exhbit quantum entanglement or EPR steering.
The new result in this paper is that a non-local hidden variable theory has been presented, and Bell inequalities obtained based on this NLHVT. These inequalities are of the Collins, Gisin, Linden, Massar Popescu (CGLMP) type.. It has been shown that in a maximally entangled state in a bipartite system clear violations of these Bell inequalities occur, thus showing that it is not posible to underpin or complete quantum theory via a non-local hidden variable theory.
It should be pointed out that the observables chosen have no obvious physical meaning. Also the experimental measurements for the NLHVT Bell inequality terms would difficult to perform. Nor would the maximally entangled state be easy to prepare. However these issues are irrelevant as all that is needed is that the observables are possible quantum observables, the state is a possible quantum state and the Bell ineqquality measurements could be carried out.
The result that a non-local hidden variable theory is also in conflict with quantum theory is of some significance since it shows that quantum theory can not be understood  in terms of hidden variable theory- as Einstein had expected. This is irrespective of whether the hidde variable theory is local or non-local.
@@ 8.7 Overall Scheme
The overall scheme can be pictured as (see [24], [16])  Here we have assumed that Bell non-local states are quantum entangled states.
@@ 9.1 Case 1 - Bell Inequality - All Conditions T
This solution is $j=k=l=m$ and here we have $\delta_{11,22}=2$ and $\delta_{12,21}=2$’  So in general wecan say that if the requirement is satisfied then $\delta_{11,22}\leq 2$ and $\delta_{12,21}\leq 2$.
Hence
$S$$\leq$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\left\{\begin{array}[]{c}\mathop{\displaystyle\sum}\limits_{j,k,l,m}P(\alpha_{j},\beta_{l}|\Omega_{A1},\Omega_{B1},\lambda)\,P(\alpha_{k},\beta_{m}|\Omega_{A2},\Omega_{B2},\lambda)\left(2\right)\\ +\mathop{\displaystyle\sum}\limits_{j,k,l,m}P(\alpha_{k},\beta_{l}|\Omega_{A2},\Omega_{B1},\lambda)\,P(\alpha_{j},\beta_{m}|\Omega_{A1},\Omega_{B2},\lambda)\left(2\right)\end{array}\right\}$
$\leq$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\;(2+2)$
$\leq$ 44 44 (54)
@@ 9.2 Case 2 - Bell Inequality - Three Conditions T, One F
Here there are four sub-cases
@@ 9.2.2 Sub-Case b - $C1=T,C2=T,C3=F,C4=T$
This is the same as for Sub_Case a. In this sub-case we have $\delta_{11,22}=1$ and $\delta_{12,21}=2$. So in general wecan say that if the requirement is satisfied then $\delta_{11,22}\leq 1$ and $\delta_{12,21}\leq 2$.
@@ 9.2.5 All Sub-Cases of Case 2
Overall, we see that if the requirement is satisfied, we can either say $\delta_{11,22}\leq 1$ and $\delta_{12,21}\leq 2$ or $\delta_{11,22}\leq 2$ and $\delta_{12,21}\leq 1$. Any of these possibilities occurs when three of $C1,C2,C3,C4$ are true and one is false.
Hence applying the first possibility
$S$$\leq$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\left\{\begin{array}[]{c}\mathop{\displaystyle\sum}\limits_{j,k,l,m}P(\alpha_{j},\beta_{l}|\Omega_{A1},\Omega_{B1},\lambda)\,P(\alpha_{k},\beta_{m}|\Omega_{A2},\Omega_{B2},\lambda)\left(1\right)\\ +\mathop{\displaystyle\sum}\limits_{j,k,l,m}P(\alpha_{k},\beta_{l}|\Omega_{A2},\Omega_{B1},\lambda)\,P(\alpha_{j},\beta_{m}|\Omega_{A1},\Omega_{B2},\lambda)\left(2\right)\end{array}\right\}$
$\leq$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\;(1+2)$
$\leq$ 33 33 (62)
The second possibility gives the same result
@@ 9.3 Case 3 - Bell Inequality - Two Conditions T, Two Conditions F
Here there are six sub-cases
@@ 9.3.1 Sub-Case a $C1=T,C2=T,C3=F,C4=F$
Since in this sub-case we have $C1=T,C2=T,C3=F,C4=F$we have $\delta_{11,22}=1$ and $\delta_{12,21}=1$. So in general we can say that if the requirement is satisfied then $\delta_{11,22}\leq 1$ and $\delta_{12,21}\leq 1$.
@@ 9.3.2 Sub-Case b $C1=T,C2=F,C3=T,C4=F$
Since in this sub-case we have $C1=T,C2=F,C3=T,C4=F$ we have $\delta_{11,22}=2$ and $\delta_{12,21}=0$. So in general we can say that if the requirement is satisfied then $\delta_{11,22}\leq 2$ and $\delta_{12,21}=0$.
@@ 9.3.3 Sub-Case c $C1=T,C2=F,C3=F,C4=T$
Since in this sub-case we have $C1=T,C2=F,C3=F,C4=T$ we have $\delta_{11,22}=1$ and $\delta_{12,21}=1$. So in general we can say that if the requirement is satisfied then $\delta_{11,22}\leq 1$ and $\delta_{12,21}\leq 1$.
@@ 9.3.4 Sub-Case d $C1=F,C2=T,C3=T,C4=F$
Since in this sub-case we have $C1=F,C2=T,C3=T,C4=F$ we have $\delta_{11,22}=1$ and $\delta_{12,21}=1$. So in general we can say that if the requirement is satisfied then $\delta_{11,22}\leq 1$ and $\delta_{12,21}\leq 1$.
@@ 9.3.5 Sub-Case e $C1=F,C2=T,C3=F,C4=T$
Since in this sub-case we have $C1=F,C2=T,C3=F,C4=T$ we have $\delta_{11,22}=0$ and $\delta_{12,21}=2$. So in general we can say that if the requirement is satisfied then $\delta_{11,22}=0$ and $\delta_{12,21}\leq 2$.
@@ 9.3.6 Sub-Case f $C1=F,C2=F,C3=T,C4=T$
Since in this sub-case we have $C1=F,C2=F,C3=T,C4=T$ we have $\delta_{11,22}=1$ and $\delta_{12,21}=1$. So in general we can say that if the requirement is satisfied then $\delta_{11,22}\leq 1$ and $\delta_{12,21}\leq 1$.
@@ 9.3.7 All Sub-Cases of Case 3
Overall, we see that if the requirement is satisfied, we can either say $\delta_{11,22}=0$ and $\delta_{12,21}\leq 2$ or $\delta_{11,22}\leq 1$ and $\delta_{12,21}\leq 1$ or $\delta_{11,22}\leq 2$ and $\delta_{12,21}=0$ Any of these possibilities occurs when two of $C1,C2,C3,C4$ are true and two are false.
Hence applying the second possibility
$S$$\leq$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\left\{\begin{array}[]{c}\mathop{\displaystyle\sum}\limits_{j,k,l,m}P(\alpha_{j},\beta_{l}|\Omega_{A1},\Omega_{B1},\lambda)\,P(\alpha_{k},\beta_{m}|\Omega_{A2},\Omega_{B2},\lambda)\left(1\right)\\ +\mathop{\displaystyle\sum}\limits_{j,k,l,m}P(\alpha_{k},\beta_{l}|\Omega_{A2},\Omega_{B1},\lambda)\,P(\alpha_{j},\beta_{m}|\Omega_{A1},\Omega_{B2},\lambda)\left(1\right)\end{array}\right\}$
$\leq$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\;(1+1)$
$\leq$ 22 22 (85)
The two other possibilities give the same result.
@@ 9.4 Case 4 Bell Inequality - One Condition T, Three Conditions F
Here there are four sub-cases.
@@ 9.4.1 Sub-Case a $C1=T,C2=F,C3=F,C4=F$
Since in this sub-case we have $C1=T,C2=F,C3=F,C4=F$ we have $\delta_{11,22}=1$ and $\delta_{12,21}=0$. So in general we can say that if the requirement is satisfied then $\delta_{11,22}\leq 1$ and $\delta_{12,21}=0$.
9.4.2 Sub-Case b $C1=T,C2=F,C3=F,C4=F$
@@
Since in this sub-case we have $C1=F,C2=T,C3=F,C4=F$ we have $\delta_{11,22}=0$ and $\delta_{12,21}=1$. So in general we can say that if the requirement is satisfied then $\delta_{11,22}=0$ and $\delta_{12,21}\leq 1$.
@@ 9.4.3 Sub-Case c $C1=F,C2=F,C3=T,C4=F$
Since in this sub-case we have $C1=F,C2=F,C3=T,C4=F$ we have $\delta_{11,22}=1$ and $\delta_{12,21}=0$. So in general we can say that if the requirement is satisfied then $\delta_{11,22}\leq 1$ and $\delta_{12,21}=0$.

### v2

@@ 4.4 Quantum Theory Test
and using the QTHY result (29) for $P(\alpha_{k},\beta_{l}|\Omega_{Aa},\Omega_{Bb},\rho)_{Q}$ in the case of the maximally entangled state we obtain
$S$ $=$ $\frac{1}{2d^{3}}\;\mathop{\displaystyle\sum}\limits_{jl}\frac{1}{\sin^{2}(\frac{\pi}{d}(j-l+\theta_{1}+\phi_{1}))}\;\delta_{j,l+\Delta_{11}}$
$+\frac{1}{2d^{3}}\;\mathop{\displaystyle\sum}\limits_{kl}\frac{1}{\sin^{2}(\frac{\pi}{d}(k-l+\theta_{2}+\phi_{1}))}\;\delta_{l,k+\Delta_{12}}$
$+\frac{1}{2d^{3}}\;\mathop{\displaystyle\sum}\limits_{km}\frac{1}{\sin^{2}(\frac{\pi}{d}(k-m+\theta_{2}+\phi_{2}))}\;\delta_{k,m+\Delta_{22}}$
$+\frac{1}{2d^{3}}\;\mathop{\displaystyle\sum}\limits_{jm}\frac{1}{\sin^{2}(\frac{\pi}{d}(j-m+\theta_{1}+\phi_{2}))}\;\delta_{m,j+\Delta_{21}}$
@@
for $S$ in the QTHY approach. Note that adding an integer multiple of $d$ to the $\Delta_{ij}$ does not alter the result.
Note that although the four sets of measurements for the result for $S$ involve one pair of the two sub-system observables at a time - ($\Omega_{A1},\Omega_{B1}$), ($\Omega_{A1},\Omega_{B1}$), ($\Omega_{A1},\Omega_{B1}$) or  ($\Omega_{A1},\Omega_{B1}$) the outcomes for the other pair of observables are disregarded. In an earlier paper [8] we pointed out that although for HVT (which involves non-quantum observables) the way the other pair of observable outcomes was disregared was unambiguous, this was not the case in quantum theory. Options included (a) not measuring the other pair  observables at all (b) measuring the other pair of observables and then discarding the results - and this could be done in two ways depending on whether the disregarded observables were measured first or second. The quantum expressions differ for (a) and (b) since the latter involve a projection operation into the space of the disregarded observables. Any of these would be a possible quantum prediction for comparison purposes. However, as it is only necessary to show HVT is inconsistent with one possible quantum prediction we choose the simplest option, namely that in each term the other observables are not measured at all.  This was the choice made by CGLMP and Eq (29) for the quantum theory expression $P(\alpha_{k},\beta_{l}|\Omega_{Aa},\Omega_{Bb},\rho)_{Q}$ is determined this way.
We could choose as a special case $\Delta_{ij}$ in the form $\Delta_{ij}=d\times\Delta_{ij}^{\ast}$ where the $\Delta_{ij}^{\ast}$ are all integers. In this case, using $\sin^{2}(\pi\Delta_{ij}^{\ast}+\eta)=(\sin(\pi\Delta_{ij}^{\ast})\cos(\eta)+\cos(\pi\Delta_{ij}^{\ast})\sin(\eta))^{2}=\sin^{2}(\eta)$ in Eq (LABEL:Eq.QThyResultS_CGLMP), we obtain for the QTHY approach
@@
We have used the values $a=1,2\quad\theta_{1}=0,\theta_{2}=1/2$ and $b=1,2\quad\phi_{1}=1/4,\phi_{2}=-1/4$ from Eqs (26), (LABEL:Eq.EigenvectorsObservB) for  the eigenvectors. This last result shows that in quantum theory the value of $S$ is independent of the $\Delta_{ij}^{\ast}$. It only depends on the integer $d$. The NLHVT Bell inequalities for $S$ do of course depend on the $\Delta_{ij}^{\ast}$.
For the choice $d=10$ we find that in QTHY $S\approx 2.52$, so this exceeds $S=2$ and $S=1$ though not $S=4$ and $S=3$ that could occur in NLHVT depending on the requirements for these cases being met. For the choice $(\Delta_{11}^{\ast},\Delta_{12}^{\ast},\Delta_{22}^{\ast},\Delta_{21}^{\ast})=(1,-2,1,-2)$ it is easily confirmed that the requirements for $S\leq 4$ are not met. However the requirements for $S\leq 3$, for $S\leq 2$ and for $S\leq 1$ are all met. . Other choices are also be made. For example with $(\Delta_{11}^{\ast},\Delta_{12}^{\ast},\Delta_{22}^{\ast},\Delta_{21}^{\ast})=(1,-3,1,-3)$ it is easily confirmed that the requirements for $S\leq 4$ are not met. However the requirements for $S\leq 3,S\leq 2$ and $S\leq 1$ are all met. In  both these  situations NLHVT predicts $S\leq 1$ this being the smallest of the three inequalities. Thus we have found cases for $d=10$ where the quantum theory prediction that $S=2.52$ is inconsistent with the NLHVT Bell inequality prediction of $S\leq 1.$
We could also choose $(\Delta_{11},\Delta_{12},\Delta_{22},\Delta_{21})=(0,1,0,0)$ as in the original CGLMP Bell inequality (18). For this choice the requirements for $S\leq 3$, for $S\leq 2$ and for $S\leq 1$ are all met. .In this case the CGLMP Bell inequality predicts $S\leq 1$. For this choice the QTHY expression for $S$ is given by
@@
For the case of $d=2$ this gives $S=3.41$. For the choice $(\Delta_{11},\Delta_{12},\Delta_{22},\Delta_{21})=(0,1,0,0)$ the requirements for $S\leq 4$ are not met, but those for $S\leq 1$ are met,.However NLHVT is inconsistent with QTHY in this choice of $(\Delta_{11},\Delta_{12},\Delta_{22},\Delta_{21})$ as it was for LHVT where in both  types of HVT the inequality $S\leq 1$ is violated in Q THY.
@@ 5 SUMMARY AND CONCLUSION
A brief review of the origins of local hidden variable theory, Bell local states and Bell inequalities has been presented. We have also outlined the basic justification for LHVT. The classification of Bell local states into three categorise based on the local hidden quantum state approach of Wiseman [16] has been explained. We outlined how the Bell inequalities provide tests for whether local hidden variable theory is in conflict with quantum theory and also which categories of Bell local states exhbit quantum entanglement or EPR steering.
The new result in this paper is that a Bell-type non-local hidden variable theory has been presented, and Bell inequalities obtained based on this NLHVT. This is based on the approach of Collins, Gisin, Linden, Massar Popescu (CGLMP) for bipartite systems, where the basic joint probability involves the outcomes for two observables for each  sub-system  In addition to showing that the well-known CHSH inequality can also be proved  for such a NLHVT, we have also considered CGLMP Bell inequalities involving the quantity $S$ that involves joint probabilities for outcomes from a pair of sub-system observables,  one in each sub-system. Sufficiency conditions for the outcome shifts have been obtained for CGLMP inequalities $S\leq 4,S\leq 3,S\leq 2$ and $S\leq 1$. It has been shown that in a maximally entangled state in a bipartite system clear quantum theory violations of these CGLMP Bell inequalities occur , thus showing that it is not posible to underpin or complete quantum theory via a CGLMP  Bell-type non-local hidden variable theory. The same sufficiency conditions also show that there is also a quantum theory violation for the CGLMP version of LHVT.
It should be pointed out that the observables chosen have no obvious physical meaning. Also the experimental measurements for the NLHVT Bell inequality terms would difficult to perform. Nor would the maximally entangled state be easy to prepare. However these issues are irrelevant as all that is needed is that the observables are possible quantum observables, the state is a possible quantum state and the Bell inequality measurements could be carried out.
The result that a Bell-type non-local hidden variable theory is also in conflict with quantum theory is of some significance since it shows that quantum theory can not be understood  in terms of Bell-type hidden variable theory. This is irrespective of whether the hidden variable theory is local or non-local.
@@ 7.7 Overall Scheme
The overall scheme can be pictured as (see [27], [23])  Here we have assumed that Quantum Theory is underpinned by Hidden Variable Theory - local or non-local and that Bell non-local states are quantum entangled states. The issue of whether or not quantum theory is underpinned by a non-local hidden variable theory is the focus of the main part of this paper.
@@ 8.1 Case 1 - Bell Inequality - All Conditions T
This solution is $j=k=l=m$ and here we have $\delta=4$
Hence applying this result for $\delta$ in Eq. (38)
$S$$\leq$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\left\{\mathop{\displaystyle\sum}\limits_{j,k,l,m}P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)\left(4\right)\,\right\}$ (63)
$\leq$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\;(4)$
$\leq$ 44 44
@@ 8.2 Case 2 - Bell Inequality - Three Conditions T, One F
Here there are four sub-cases. In all of thes sub-cases we have $\delta=3.$
@@ 8.2.5 All Sub-Cases of Case 2
Overall, we see that if the requirement is satisfied, we can say $\delta=3.$
Hence applying this result for $\delta$ in Eq. (38)
$S$$\leq$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\left\{\mathop{\displaystyle\sum}\limits_{j,k,l,m}P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)\left(3\right)\,\right\}$ (68)
$\leq$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\;(3)$
$\leq$ 33 33
@@ 8.3 Case 3 - Bell Inequality - Two Conditions T, Two Conditions F
Here there are six sub-cases. In all of these sub-cases $\delta=2$.
@@ 8.3.7 All Sub-Cases of Case 3
Overall, we see that if the requirement is satisfied, we can say $\delta=2$.
Hence applying this value of $\delta$ in Eq. (38)
$S$$\leq$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\left\{\mathop{\displaystyle\sum}\limits_{j,k,l,m}P(\alpha_{j},\alpha_{k},\beta_{l},\beta_{m}|\Omega_{A1},\Omega_{A2},\Omega_{B1},\Omega_{B2};\lambda)\left(2\right)\,\right\}$ (88)
$\leq$ $\mathop{\displaystyle\sum}\limits_{\lambda}P(\lambda|c)\;(2)$
$\leq$ 22 22
@@ 8.4 Case 4 Bell Inequality - One Condition T, Three Conditions F
Here there are four sub-cases. In all these  sub-cases  $\delta=1$.
@@ 8.4.1 Sub-Case a $C1=T,C2=F,C3=F,C4=F$
8.4.2 Sub-Case b $C1=F,C2=T,C3=F,C4=F$